OpenStudy (anonymous):

Simplify sqrt[(169)(x^6)(y^8)]

jimthompson5910 (jim_thompson5910):

\[\large\sqrt{169x^6y^8}\] \[\large\sqrt{(13)^2(x^3)^2(y^4)^2}\] \[\large\sqrt{(13)^2}\sqrt{(x^3)^2}\sqrt{(y^4)^2}\] \[\large13x^3y^4\] So \[\large\sqrt{169x^6y^8}=13x^3y^4\]

jimthompson5910 (jim_thompson5910):

This is of course assuming all variables are nonnegative.

OpenStudy (anonymous):

well just x doesnt matter for y

jimthompson5910 (jim_thompson5910):

good point jahtoday

OpenStudy (anonymous):

you guys are awsome that helped a bunch!